OpenStudy (anonymous):

Which value is a solution for the equation cot (x/2) = 1 ? A. 5π/2 B. 7π/4 C. π/4 D. 3π/4

OpenStudy (anonymous):

I know that x = 1/2 (4π n+π) ..... but what does that mean?

OpenStudy (mathstudent55):

\( \cot x = \dfrac{1}{\tan x} \) If \( \cot \frac{x}{2} = 1\) then \( \dfrac{1}{\tan \frac{x}2{}} = 1 \) and \( \tan \frac{x}{2} = 1 \) Where is tan x = 1?

OpenStudy (anonymous):

x = 1/4 (4 π n+π) So does that mean it is C. ?

OpenStudy (mathstudent55):

\( \tan \theta = 1\) at \( \theta = \dfrac{\pi}{4} \) \(\tan \frac{x}{2} = 1 \) \( \dfrac{x}{2} = \dfrac{\pi}{4} \) \( x = \dfrac{\pi}{2} \)

OpenStudy (anonymous):

Nvm. Cant be! Must be A.

OpenStudy (mathstudent55):

Since the tangent function is a cyclical function, any multiple of \(2 \pi \) you add to it is also a solution. x = 1/2 (4π n+π) = \( 2 \pi n + \dfrac{\pi}{2} \), for all integers, n. There you see the pi/2 of our solution plust the \(2 \pi n \) part which is all the other solutions. All you need to do is see which choice is a multiple of \(2 \pi \) added to \( \dfrac{\pi}{2} \)

OpenStudy (mathstudent55):

Correct. It's A.